solve by the elimination method, what is the ordered pair?
3a+2b=-25
-3a+b=28

Answer by jim_thompson5910(35256)

Hint: add the equations to get


3a+2b=-25
-3a+b=28
-----------
0a+3b = 3


So 3b = 3 which means b = 1 (divide both sides by 3)


Use b = 1 to find the value of 'a'.